### v3

# PerfTracker: Online Performance Diagnosis for Large Model Training

Abstract.
Troubleshooting performance problems of large model training (LMT) is immensely challenging, due to unprecedented scales of modern GPU clusters, the complexity of software-hardware interactions, and the data intensity of the training process. Existing troubleshooting approaches designed for traditional distributed systems or datacenter networks fall short and can hardly apply to real-world training systems. In this paper, we present PerfTracker, the first online troubleshooting system utilizing fine-grained profiling, to diagnose performance issues of large-scale model training in production. PerfTracker can diagnose performance issues rooted in both hardware (e.g., GPUs and their interconnects) and software (e.g., Python functions and GPU operations). It scales to LMT on modern GPU clusters. PerfTracker effectively summarizes runtime behavior patterns of fine-grained LMT functions via online profiling, and leverages differential observability to localize the root cause with minimal production impact. PerfTracker has been deployed as a production service for large-scale GPU clusters of $O(10,000)$ GPUs (product homepage https://help.aliyun.com/zh/pai/user-guide/perftracker-online-performance-analysis-diagnostic-tool). It has been used to diagnose a variety of difficult performance issues.
copyright: nonedoi: isbn: price:
1. Introduction
Troubleshooting performance problems of Large Model Training (LMT) has become a grand challenge driven by unprecedented increases in AI model sizes and GPU-based training infrastructures. LMT undertakes complex hardware-software interactions; unexpected behavior of hardware components (e.g., GPUs and their interconnects) to software code and configurations (e.g., data loaders, communication, and PyTorch), could lead to performance issues. With rapid scaling of LMT, occurrences of performance problems inevitably increase accordingly, resulting in significant wastage of compute resources and production incidents (Jiang et al., 2024; Qian et al., 2024; Xiong et al., 2024).
Unfortunately, traditional troubleshooting techniques, designed for distributed systems (Zhao et al., 2014; Barham et al., 2004; Terai et al., 2014; Harsh et al., 2023) and data center networks (Liu et al., 2024; Tan et al., 2019; Arzani et al., 2018; Roy et al., 2017), fall short and can hardly help today’s model training systems. First, all existing approaches focus on CPU-based systems and do not work with GPU-centric, heterogeneous systems. Modern LMT systems are orders of magnitude larger in scale than traditional distributed systems and networks in terms of parallel computation. The scale and parallelism impair existing approaches that rely on fine-grained traces and profiling data. For example, an LMT of 10,000 GPUs typically makes $4\times 10^{9}$ function executions per second, which produces terabytes of profiling data. Active measurement (e.g., probing system components (Liu et al., 2024; Tan et al., 2019)) is not an option due to the interference with production workloads. Moreover, most existing approaches focus only on software defects or only on hardware infrastructures; they cannot reason about their complex interactions.
In practice, troubleshooting performance issues of production LMT relies on data from either online monitoring or offline profiling. Unfortunately, neither is ideal and engineers often face difficult trade-offs between observability and cost. Online monitoring has to be coarse-grained to avoid production impact, while offline profiling loses timeliness and requires a reproduction of the observed problem in a small test environment (which is not always possible).
Most online monitors use vendor tools to collect performance counters of hardware components (e.g., GPUs (dcg, 2024), CPU (pcm, 2024), CPU-GPU (dyn, 2024), and NICs (pcm, 2024; mst, 2024)) during LMT. To avoid production impacts, the sampling can only be coarse-grained, typically in seconds. Similarly, monitors on function executions can only record coarse-grained runtime information like forward phase, backward phase, and CUDA kernels (meg, 2019; Jiang et al., 2024). However, coarse-grained observability cannot capture fine-grained operations in LMT, which is critical for performance analysis. In our experience, online monitors can only detect coarse-grained symptoms (e.g., performance degradation in forward/backward phases or in communication modules), but cannot pinpoint fine-grained root causes (e.g., which function or network link caused the degradation).
Offline profiling (e.g., Torch Profiler (tor, 2024), Nsight System (nsy, 2024), and Nsight Compute (nsi, 2024)) records fine-grained operations during LMT execution, including both function executions (e.g., Python call stacks, CUDA executions) and high-frequency hardware sampling (e.g., at 200k Hz). Based on these profiling data, a troubleshooting system can detect subtle anomalies of program behaviors. But, given the sheer volume of profiling data across the stack, analyzing data of all workers at runtime cannot be done at scale. To leverage profiling, engineers must manually reproduce the issues in a small-scale LMT offline, which is time-consuming and not always possible, leading to prolonged recovery time.
We aim to develop a troubleshooting system to effectively localize fine-grained root causes for production issues. The key challenge is to offer fine-grained observability as in offline profiling, while being lightweight for online use cases.
Contributions. We present PerfTracker, the first online performance troubleshooting system that can use fine-grained profiling data for large-scale model training in production. PerfTracker pinpoints root causes of performance issues in both hardware and software components. It scales to modern AI models on large-scale GPU infrastructure— identifying abnormal function executions for a 3,072-GPU LMT only takes 3 minutes (7 minutes for a 1,000,000-GPU LMT). PerfTracker supports all training frameworks using PyTorch. To use PerfTracker, one only needs to import it in the Python source code for LMT without other modifications.
The key insight of PerfTracker is to develop highly efficient differential observability of LMT. PerfTracker does not attempt to directly analyze raw profiling data of all the workers (e.g., compare all function events between all the workers). Instead, PerfTracker summarizes the runtime behavior pattern of causally related LMT functions on the critical path of the observed performance anomaly. We show that it is possible to concisely summarize each function’s runtime behavior patterns using statistical metrics (e.g., execution duration, hardware resource utilization, etc.), without loss of diagnosability. Since the size of runtime behavior patterns is orders of magnitudes smaller than raw profiling data, the abnormal behavior can be efficiently identified by pattern analysis, leveraging the asymmetric architecture of model training.
We have developed PerfTracker as a production service in our LMT infrastructure with $O(10,000)$ GPUs and deployed it for three months (product homepage https://help.aliyun.com/zh/pai/user-guide/perftracker-online-performance-analysis-diagnostic-tool). PerfTracker has been helping our customers to troubleshoot various LMT performance issues which can hardly be addressed by existing practices or state-of-the-art techniques, including hardware and code-level issues.
Summary. This paper makes the following contributions:
  • We present PerfTracker as the first profiling-based online performance troubleshooting system for LMT in production.
  • We describe the technique of summarizing runtime behavior patterns of LMT functions, which enables highly efficient differential observability for LMT.
  • We share our production results of PerfTracker.
Ethics. This work does not raise any ethical issues.
2. Background and Motivation
2.1. Performance Issues in LMT
Large model training (LMT) requires the cooperation of multiple hardware and software modules (Figure 1). The hardware typically includes GPUs (NVI, 2024a, b), CPUs, DRAM, PCIe, NVLinks (NVL, 2024), NICs (CX6, 2024; CX7, 2024), and inter-host networks (Greenberg et al., 2009; Qian et al., 2024). Software modules include GPU operators, communication libraries like NCCL (NCC, 2023), training frameworks (e.g., Megatron (meg, 2019), Nemo (nem, 2024)), containers, etc.. Malfunctions in any of the above components can lead to performance issues, such as throughput decrease, fluctuations, or training blockage. In this paper we do not focus on LMT crashes, which can be explicitly identified by error messages.
Figure 1. The full stack of large model training.
Figure 2. An overview of LMT performance issues.
Figure 2 breaks down LMT performance issues in our production GPU cluster with $O(10,000)$ GPUs in the past nine months. These issues manifest as slow computation and communication, but with diverse root causes—44.4% of them are caused by hardware issues, and 48.2% are software issues.
Computation slowdowns can be caused by hardware (e.g., GPU throttling), deficient user code, misconfigurations, or incompatible libraries. For example, in our customers’ models, some training code frequently transfers data between CPUs and GPUs, introduces excessive synchronization, uses outdated PyTorch versions, poorly manages memory, or contains custom functions that run unexpectedly long.
Communication slowdowns can stem from physical networks, including NVLink, PCIe, NICs, and switches with optical modules. Although major communication throughput reduction is easy to detect, pinpointing the exact machines or hardware components is much more challenging. The difficulty is because most communications in LMTs are collective operations—a slowdown in any network link can proportionally reduce the communication throughput of many workers simultaneously. Besides, bugs in communication libraries, RDMA, and workload managers (e.g., Slurm (slu, 2024)) can also lead to communication slowdowns.
Furthermore, system services used by LMT can lead to performance issues as well. For example, storage services can affect the performance of data loading or checkpointing during LMT. Management services like load tests and monitoring are co-located on each host; unexpected behavior can lead to resource contention, slowing down the LMT.
2.2. Common Practices
Before PerfTracker, we relied on existing tools or techniques to troubleshoot performance issues of LMT (Table 1). However, none of them meets our needs in production.
Monitoring. We deployed online monitors on hardware performance counters of GPUs, PCIe, NVLink, and the inter-host network with second-level sample rates, using existing tools (e.g., DCGM (dcg, 2024), Dynolog (dyn, 2024)) and state-of-the-art techniques (e.g., Minder (Deng et al., 2024) and MegaScale (Jiang et al., 2024)); we also deployed monitors to trace forward and backward time and CUDA execution time. However, as shown in Figure 2, only 29.6% of performance issues can be detected.
The key reason is that online monitors are coarse-grained and are unaware of LMT behavior, creating a fundamental gap between the observation and the root cause. First, performance issues caused by configurations or code are not visible via hardware monitoring. Second, most warnings from monitors do not necessarily indicate performance issues in LMT; many abnormal indicators can also be a result of temporarily high pressure on hardware (e.g., excessive CNPs) or correctable errors (e.g., XID 63,64, and 92), which do not always degrade the end-to-end performance of LMT. Third, even when performance issues are caused by problems covered by our monitors (e.g., hardware issues), they may not be detectable due to coarse-grained online monitoring.
In practice, engineers usually analyze the code behavior of LMT in conjunction with hardware throughput and warning information, which have not been automated online.
Profiling. Profilers can analyze fine-grained behavior of LMT and are used as a supplement to online monitors. Profilers like Nsight, Torch Profiler, and Python Profiler record all activities in LMT execution with microsecond-level hardware sampling, which provide sufficient information to identify the root cause of sophisticated performance issues.
However, given their significant overhead (e.g., Torch Profiler generates 100+ MB profiling data for one worker per second), profiling can only be done in a small testbed. Engineers must first reproduce the observed performance issues in production in the testbed before using a profiler. However, reproducing production issues is not always possible or can be done timely. Besides problems that only manifest at a large scale, a practical problem is that user code is often confidential and cannot be disclosed to the GPU cluster providers. In practice, engineers often have to reproduce production issues using similar open-source models, which further reduces the chances of reproduction. Although we invested tremendous efforts to reproduce LMT performance issues reported by our customers, there are still 7.4% of performance issues that cannot be reproduced. For the ones that we eventually reproduced, it takes days to months of effort.
Table 1. Performance diagnosis tools for LMT
Usage Tool Diagnostic Information
GPU/DRAM/ PCIe/NVLink NIC Python Events Kernels
Offline Nsight Systems (nsy, 2024) $<$200kHz $<$1kHz
Torch Profiler (tor, 2024)
Cupti (cup, 2024)
Online DCGM (dcg, 2024) 1Hz
MegaScale (Jiang et al., 2024) $<$1kHz
Dynolog (dyn, 2024) 0.1Hz 0.1kHz
Minder (Deng et al., 2024) $<$1Hz $<$1Hz
Online PerfTracker $<$200kHz $<$1kHz
2.3. Goals and Challenges
Our goal is to significantly improve both efficiency and effectiveness of LMT performance troubleshooting in production at scale, with (1) fine-grained observability: collecting all function execution events during LMT, with fine-granularity hardware sampling, (2) efficiency and scalability: as fine-grained observability leads to large volumes of data, online troubleshooting must be efficient and scalable to large GPU clusters; (3) low overhead: introducing no performance impact on the routine LMT training, and minimizing the requirements for computation power, network bandwidth, and storage. We address the following challenges:
Challenge 1: Overhead of fine-grained observability.
  • $\bullet$
    Profiling generates a huge amount of data. The size of profiling data of a single worker, including application-level tracing (e.g., the timeline of function execution) and hardware metrics (e.g., using nsys (nsy, 2024) for GPU, NVLink, and PCIe, in 10k Hz), is typically 100 MB per second. Suppose an LMT requires 10,000 GPUs, a profiler can generate $\sim$1 TB data per second, which is untenable.
  • $\bullet$
    Profiling may reduce LMT performance. Profiling may reduce LMT training throughput (§6.3). Most customers cannot afford periodic profiling for their LMT jobs.
Challenge 2: Needles in the haystack.
  • $\bullet$
    Extracting important information. Fine-grained profiling data far exceeds the volume that can be efficiently analyzed in real time (e.g., 1 TB/s for a 10,000-GPUs LMT). In fact, it is even hard to aggregate all the data; so, algorithms that assume seeing all the data are hardly practical. Dropping or downsampling is not ideal, as it may reduce important details and thus impair the diagnosability.
  • $\bullet$
    Avoid expensive coordination. In production, hosts are not always perfectly synchronized in clock time. With protocols like Network Time Protocol (NTP) (Mills, 1991), the error could be $\sim$10 milliseconds. Algorithms that compare timestamps of events between workers would not work—a lot of function executions are on the millisecond or even microsecond scale. A decentralized algorithm is needed.
3. Main Idea
The key idea of PerfTracker is to develop highly efficient differential observability of LMT functions, leveraging homogeneous architectures of LMT. We use the term “function” to refer to any procedure in LMT, including Python functions, GPU/CPU kernel functions, memory operations, etc.. Compared to traditional distributed systems, LMT function executions (examples in Appendix A) exhibit two significant characteristics.
  • $\bullet$
    Within a single worker, most low-level functions are executed repeatedly. This is because training involves many identical iterations, and models are typically composed of repetitive submodules (e.g., transformer blocks).
  • $\bullet$
    Runtime behaviors of functions are highly identical across workers, because modern parallelisms (e.g., at data, pipeline, tensor, and expert levels) distribute workloads evenly with frequent synchronization operations.
Therefore, most performance issues can be diagnosed by observing abnormal function runtime behaviors in comparison to all other function executions. We elaborate on this idea using a representative example of an LMT performance issue caused by ring communication functions.
Ring communication in LMT (like NCCL (NCC, 2023) AllReduce and AllGather) connects workers in ring-like topologies for direct data exchange with only their adjacent neighbors. A ring topology utilizes both NVLinks (intra-host) and GPU-NIC (inter-host) links. In this example, a NCCL AllReduce group with 32 GPUs is distributed on 4 hosts, each hosting 8 GPUs, with every pair of GPUs sharing two bonded NICs. A malfunctioning NIC downgrades a bond by 50% and creates a performance issue for the NCCL AllReduce function. Since multiple rings typically share the intra-host network (e.g., NVLink) but use different NICs for inter-host communication, we focus on each worker’s GPU-NIC throughput during NCCL AllReduce (Figure 3), to distinguish the communication behaviors of different rings and different workers.
After downgrading a network bond, all workers’ GPU-NIC throughputs are categorized into three different patterns.
  • $\bullet$
    For workers not included in a ring using this downgraded bond (green nodes in Figure 4), the throughput pattern (Figure 55(a)) is identical to the healthy state (Figure 3).
  • $\bullet$
    For workers in a ring using the downgraded bond but not directly connecting to this bond (blue nodes in Fig. 4), the average throughput is lower (Fig. 55(b)) with fluctuations.
  • $\bullet$
    The worker that directly connects to the downgraded bond (red node in Figure 4) also has low average throughput (Figure 55(c)) but is more stable than Figure 55(b).
Figure 3. PCIe bandwidth utilization (from GPU to NIC) during a ring communication without network issues.
Figure 4. The ring topology with and without a problematic link. We illustrate a link with only six workers for simplicity. There are 32 in our experiment.
(a) GPU-NIC links at maximal throughput.
(b) GPU-NIC links of lower throughput and high fluctuation.
(c) GPU-NIC link of lower and stable throughput (slow link).
Figure 5. GPU-NIC throughput pattern. Left figures are the bandwidth utilizations during the execution of a ring communication function, and right figures show enlarged views of local areas for better observation.
Figure 6. Overview of the PerfTracker system.
To explain the low average throughput of Figures 55(b) and 55(c) compared with Figure 55(a): during the ring communication, the NCCL communication library constructs multiple rings, each using different NICs, to link all workers head-to-tail. If a ring contains no slow link, all of its links operate at maximal throughput. For rings including a slow link, the average throughput of all workers included in this ring drops.
To explain the fluctuation in Figure 55(b): ring communication always transfers data in small chunks one by one. At each stage, each worker transfers a chunk to the next worker, and does some computation based on the chunk it received. With a slow link, though normal links in the ring can still reach maximal throughput, they have to wait for the slow link after each stage of chunk transmission. So their throughput fluctuates between zero and the maximal, with a high standard deviation. The slow link does not need to wait, so its throughput remains low with a small standard deviation.
Insights. The above example demonstrates three important opportunities: (1) performance issues can be observed by profiling the behavior of function executions; (2) we can troubleshoot performance issues using differential observability that localizes the offending function executions with abnormal behavior (e.g., low average GPU-NIC throughput without fluctuation); and (3) we do not need to analyze fine-grained raw observability data of all the functions; instead, we only need to summarize their runtime behavior patterns. In the ring communication example, each worker only needs to provide two numbers (mean and standard deviation) as the pattern of the ring communication function. In summary:
Most (if not all) LMT performance issues can be diagnosed by highly efficient differential observability on runtime behavior patterns of individual LMT functions.
4. PerfTracker Design
PerfTracker is an online performance troubleshooting system for large-scale model training (LMT). It embodies the insights discussed in §3 in a production service for our customers who run LMT on our GPU clusters. PerfTracker employs a highly efficient approach to detecting performance degradation without continuous online profiling, introducing little overhead on LMTs. When a performance degradation is detected, PerfTracker automatically troubleshoots the performance issue. PerfTracker uses online profiling to collect fine-grained information, but it does not try to aggregate raw profile data of every function; instead, PerfTracker summarizes the runtime behavior patterns of each function at every worker node and aggregates only the concise behavior patterns. The diagnosis is performed on the aggregated behavior patterns based on the differential observability principle.
Figure 6 gives an overview of PerfTracker. PerfTracker has three main components: (1) detecting performance degradation of LMT to trigger online profiling, (2) summarizing runtime behavior patterns of each function from raw profiling data, and (3) a centralized localization algorithm that pinpoints the root-cause function based on the behavior patterns. Both (1) and (2) are deployed per worker, while (3) is global.
Usage. To deploy PerfTracker, GPU cluster providers do not need access to the user code. The user only needs to import PerfTracker in the LMT code (import PerfTracker); no other code change is needed. Behind the line of import, PerfTracker initializes the detector of performance degradation, registers handler functions to start profiling, and creates a daemon process for summarizing and uploading behavior patterns. Figure 7 illustrates the output of PerfTracker. As a function-centric approach, PerfTracker pinpoints which functions on which workers are executed abnormally, and describes how they behave differently from expectation or from other workers. In Figure 7, all workers encounter slow socket receive in data loading, indicating slow storage I/O. Worker 7 is slow in collective communication, indicating a network degradation on its connected link. In addition, worker 0,1,2, and 3 execute GEMM (a CUDA kernel function (gem, 2024)) slowly because of not fully utilizing GPUs, indicating the problem of GPU throttling.
Figure 7. An example of PerfTracker output.
4.1. Detecting Performance Degradation
Profiling an LMT task online introduces overhead. To minimize the overhead while collecting profiling data timely when performance issues occur, PerfTracker automatically detects performance degradation and then makes profiling.
For an LMT task, the time spent on one training iteration is the key performance metric. PerfTracker devises a solution to monitor the training time of each iteration without accessing user code, and then triggers online profiling if the iteration time is abnormal. Note that approaches that require accessing user data (such as logs) do not work with our usage model.
Figure 8. PerfTracker’s performance degradation detection.
Indicators of iteration time. A PyTorch training iteration always involves several dataloader.next() calls, followed by several optimizer.step() calls (the number depends on training parameters like pipeline parallelism). In this sequence of events, the duration from the first dataloader.next() to the last optimizer.step() is regarded as the duration of a complete training iteration.
When it is import-ed, PerfTracker wraps the two functions by adding time counters. Basically, PerfTracker replaces the two PyTorch functions with the wrapped versions. Since the two functions are Python functions (which are not compiled before LMT executions), the replacement is done at runtime.
Iteration detection. During an LMT, PerfTracker continuously records event sequences of dataloader.next() and optimizer.step(). After detecting $M$ (=10 in practice) identical sequences starting with dataloader.next() and ending with optimizer.step(), this sequence is defined as the training iteration sequence.
Performance degradation detection. PerfTracker continuously monitors incoming events of dataloader.next() and optimizer.step() with the identified training iteration sequence. Each time PerfTracker successfully matches a complete training iteration sequence, it records the duration of that iteration. PerfTracker considers that performance degradation has occurred in the following two situations: (1) The average duration of the recent $N$ (=50 in practice) iterations exceeds the recent shortest iteration time by more than 5%. (2) The current training iteration sequence has not yet been fully matched, but the time elapsed since the last event received is at least 5×5$\times$5 × the average iteration duration (indicating the training is blocked). If PerfTracker fails to match a training iteration after $K$ (=200 in practice) consecutive event receptions, it goes back to the previous iteration detection phase to redetect the training iteration sequence. While this does not often happen in common LMT programs, it significantly enhances the robustness of PerfTracker when users implement special functionalities in their code, to make sure the algorithm always works. Figure 8 illustrates the procedure.
Profiling data generation. In PerfTracker, each LMT worker is connected to a PerfTracker daemon. Upon detection of performance degradation, all daemons are notified via TCP sockets. Each daemon sends a signal to the corresponding worker. The worker receives the signal and starts profiling via a handler function (registered when PerfTracker is imported). This ensures that profiling is executed in the main thread of LMT (or else some APIs may fail, like CUPTI in Torch Profiler).
A profiling session lasts for 20 seconds by default (which is configurable). We apply Torch Profiler to obtain function execution events for Python functions, CPU operations, memory operations, and CUDA kernels. We utilize nsys to sample hardware metrics at 10 kHz, such as GPU, DRAM, NVLink, PCIe, and the network. The profiling data is then collected by the PerfTracker daemon of each worker.
4.2. Summarizing Behavior Patterns
PerfTracker summarizes behavior patterns of each function at the worker level. Note that PerfTracker only concerns functions on the critical path of LMT (Figure 9)—only these functions contribute to the end-to-end performance, despite the huge number of function executions in an LMT.
Figure 9. The critical path of LMT performance.
PerfTracker’s solution is (1) finding the function execution events on the critical path, including GPU computation kernels, collective communication functions, memory operations, Python functions, and all other functions executed in LMT; and (2) clustering all execution events of each function (for Python functions, the entire call stack must be identical to be considered the same function), then defining several patterns to summarize the behavior of each function.
Critical path. As shown in Figure 9, different types of functions in LMT have different priorities: GPU computation kernel functions have the highest priority; memory operations (e.g., malloc, memcpy) have second priority; collective communication kernel functions (e.g., NCCL AllReduce) have third priority; and Python functions have fourth priority. A function’s execution (or a subinterval of it) is on the LMT’s critical path if and only if no higher-priority function is executing during that time. For Python functions, we additionally require that the function belongs to the training thread of the LMT (i.e., not a thread initiated by _bootstrap), and that no child functions called by this function are executing.
The rationale for such a definition is that, in a well-optimized LMT, the majority of time should be spent on GPU computations and their associated memory operations. Collective communications and Python functions should try to overlap with GPU computations as much as possible. In performance diagnosis, we only focus on the non-overlapping parts.
Runtime behavior patterns. For each function $f$ running on worker $w$, we define its pattern of behavior $P_{f,w}$ as a 3-dimensional vector:
(1) $P_{f,w}=(\beta_{f,w},\mu_{f,w},\sigma_{f,w})$
$\beta_{f,w}$ represents the percentage of the profiling window that function $f$ spends time on the critical path of worker $w$.
(2) $\beta_{f,w}=\frac{\int_{t\in T}{C(f,w,t)}}{|T|}$
where $T$ is the duration of the profiling window (e.g., 20 seconds in §4.1), and
(3) $C(f,w,t)=\begin{cases}1&\text{$f$ is on $w$'s critical path at time $t$},\\ 0&\text{Otherwise }.\end{cases}$
Base on the definition, we have $\beta_{f,w}\in[0,1]$.
$\mu_{f,w}$ represents the average hardware resource utilization during all executions of function $f$ on worker $w$. In this context, “resource” refers to hardware resources that determine performance of function $f$. For example, GPU computation kernels correspond to GPU SM utilization, Python functions to CPU utilization, intra-worker collective communication to NVLink utilization, and inter-worker collective communication corresponds to the bandwidth utilization between GPU and NIC, due to the significantly lower NIC bandwidth compared to intra-worker NVLink bandwidth.
(4) $\mu_{f,w}=\frac{\sum_{e\in E_{f,w}}{|L(e)|avg_{U}(L(e))}}{\sum_{e\in E_{f}}{|L(e)|}}$
where $E_{f,w}$ represents all execution events of function $f$ on worker $w$ in the time window, so $e$ represents each one of them. $L(e)$ represents the critical execution duration of $e$. $avg_{U}(L(e))\in[0,1]$ represents the average resource utilization during the critical execution duration. We have $\mu_{f,w}\in[0,1]$.
Figure 10. Critical duration of a collective communication function. A worker enters the communication early often need to wait for others, leaving several empty intervals of resource usage (noise). We only consider the critical duration in $\mu_{f,w}$ computation.
The definition of critical execution duration, $L(e)=[l_{c},r_{c}]$, is not trivial, and is a subinterval of $f$’s function execution duration $[l,r]$, as shown in Figure 10. This is because in practical LMT tasks, particularly within collective communication functions, numerous synchronization operations are involved. Workers that enter the collective communication kernel earlier usually perform a part of communication and then wait for others. Consequently, the throughput of network transmission are usually not continuous. The average resource utilization over the entire function execution does not accurately reflect the real communication performance. Therefore, $L(e)$ is defined as the longest interval with intense resource usage among the whole function execution duration. We apply Algorithm 1 to identify $L_{e}$ with discrete sampling of hardware resource usage.
Algorithm 1 Finding the critical execution duration
0:  Resource utilization samples $U=[U_{1},U_{2},\ldots,U_{n}]$ of function $f$’s execution duration $[l,r]$, where $U_{i}\in[0,1]$ for $\forall i$
0:  Critical execution duration $[l_{c},r_{c}]$
1:  Total resource utilization $S=\sum_{i=1}^{n}U_{i}$
2:  Initialize search bounds: $g_{\text{left}}\leftarrow 0$, $g_{\text{right}}\leftarrow n$
3:  while $g_{\text{left}}\leq g_{\text{right}}$ do
4:     $g\leftarrow$ midpoint of $[g_{\text{left}},g_{\text{right}}]$
5:     if there exists a subinterval $[l^{\prime},r^{\prime}]$ satisfying:
  1. (1)
    $\sum_{i=l^{\prime}}^{r^{\prime}}U_{i}\geq 0.8\times S$
  2. (2)
    No more than $g$ consecutive zeros in $U_{l^{\prime}},\ldots,U_{r^{\prime}}$
6:      Update $g_{\max}\leftarrow g$, $[l_{c},r_{c}]\leftarrow[l^{\prime},r^{\prime}]$
7:      $g_{\text{right}}\leftarrow g-1$
8:     else
9:      $g_{\text{left}}\leftarrow g+1$
10:  end while
11:  return  Subinterval $[l_{c},r_{c}]$
$\sigma_{f,w}$ represents the standard deviation of resource usage during all function executions of function $f$ on worker $w$, and its definition is similar to $\mu_{f,w}$:
(5) $\sigma_{f,w}=\frac{\sum_{e\in E_{f,w}}{|L(e)|std_{U}(L(e))}}{\sum_{e\in E_{f}}{|L(e)|}}$
and we also have $\sigma_{f,w}\in[0,1]$.
Data size of patterns.
(a) Raw profiling data (3GB)
(b) Runtime behavior patterns (30KB)
Figure 11. The data size of a single worker’s runtime behavior patterns compared with its raw profiling data.
Figure $1$(b) shows the size of runtime behavior patterns (30 KB) is $10^{5}\times$ smaller than the raw profiling data ($\sim$3 GB). Patterns of Python functions contributes to the majority of the data volume, because PerfTracker records the full call stack of each function, which is long (sometimes the call stack of a Python function includes 1,000 letters).
4.3. Localization
PerfTracker localizes the root causes based on the runtime behavior patterns in §4.2. In LMT execution, all performance issues can be categorized into two types: (1) a common problem of all LMT workers, like hardware misconfigurations and low-efficiency code implementation, or (2) a special problem on only a part of workers, like hardware issues. Correspondingly, given function $f$ executed on worker $w$, PerfTracker defines two types of distances, (1) Distance from expectation and (2) Differential distance, to measure its abnormality.
Distance from expectation. The first is the distance from expectation, $D_{f,w}$. In a well-optimized LMT, we have an expected range for the runtime behavior patterns of most functions (e.g., a Python function $f$ are expected to have a close-to-zero value of $\beta_{f}$ because LMT should not be bottlenecked by CPU). Any function with a runtime behavior pattern far from this range can be considered abnormal. When there are a lot of workers with large distance from expectation, it indicates common issues in the LMT, such as low hardware configurations across the cluster or deficient implementations of certain functions deployed on all workers.
To define the distance from expectation, we first need an expected range of patterns for function $f$, $R_{f}$:
(6) $R_{f}=[\beta_{f}^{l},\beta_{f}^{r}]\times[\mu_{f}^{l},\mu_{f}^{r}]\times[\sigma_{f}^{l},\sigma_{f}^{r}]$
$R_{f}$ is assigned based on our production experience. For example, if $f$ is a Python function, we have $R_{f}=[0,0.01]\times[0,1]\times[0,1]$, meaning that we do not expect an LMT to be bottlenecked by a Python function for ¿1% time. If $f$ is a collective communication function, we have $R_{f}=[0,0.3]\times[0,1]\times[0,1]$. As for GPU computation kernel functions, we have $R_{f}=[0,1]\times[0,1]\times[0,1]$, i.e., a GPU computation kernel function is never considered unexpected.
Then we can define $D_{f,w}$ as the minimal Manhattan distance from $P_{f,w}$ to $R_{f}$:
(7) $D_{f,w}=\min_{p\in R_{f}}{d_{\text{Manhattan}}{(P_{f,w},p)}}$
Differential distance. The second is the distance to other workers, $\Delta_{f,w}$. In a large-scale LMT, the behaviors of different workers should be highly consistent. If some workers exhibit significantly different function behaviors compared to the same functions executed on other workers, it often suggests that these workers are performing special operations that slow down overall training throughput (e.g., different in $\beta$-dimension), or that the hardware used by these workers for the function has performance issues that need to be addressed (e.g., different in $\mu$ or $\sigma$-dmension).
To make comparison between workers, first we make a max normalization to $P_{f,w}$:
(8) $\hat{P_{f,w}}=(\frac{\beta_{f,w}}{\max_{w\in W}{\beta_{f,w}}},\frac{\mu_{f,w}}{\max_{w\in W}{\mu_{f,w}}},\frac{\sigma_{f,w}}{\max_{w\in W}{\sigma_{f,w}}})$
Then we define $\Delta_{f,w}$ to indicate how many workers have different patterns of $f$’s execution from worker $w$:
(9) $\Delta_{f,w}=\frac{\sum_{w^{\prime}\in W_{N}}{I(\hat{P_{f,w}},\hat{P_{f,w^{\prime}}})}}{N}$
where $W_{N}$ is a subset of $N$ workers randomly sampled from the set of all workers $W$. In practical we set $N=\min{(100,|W|)}$. $I$ is a function indicating whether the distance between two vectors are below $\delta$:
(10) $I(x,y)=\begin{cases}0&\text{if }d_{\text{Manhattan}}(x,y)<\delta,\\ 1&\text{otherwise.}\end{cases}$
In practical, we set $\delta$=0.4 based on our experience.
Note that our definition of $\Delta_{f,w}$ is not based on how far a worker is from other workers (e.g., simply average the distance to other workers), but how unique a worker’s behavior is. This is because the pattern $P_{f,w}$ is in three dimensions, and each dimension has different physical meaning. A larger Manhattan distance does not mean a worker’s behavior is more abnormal (§3 shows a classical example).
Localizing abnormal function executions. Based on the two distances, a function $f$ executed on worker $w$ is regarded abnormal if:
(11) $\beta_{f,w}>0.01\ \land\ (D_{f,w}>0\ \lor\ \Delta_{f,w}>M_{f}+kMAD_{f})$
where $M_{f}=median\{\Delta_{f,w^{\prime}}|w^{\prime}\in W\}$ is the median value of distance to peers, and $MAD_{f}=median\{|\Delta_{f,w^{\prime}}-M_{f}|\ |w^{\prime}\in W\}$ is Median Absolute Deviation (mad, 2024), a robust measure of statistical dispersion. $k$ is set to 5.
$\beta_{f,w}$¿0.01 means the function execution contributes at least 1% to end-to-end performance. Otherwise, we do not need to care about this function since the end-to-end performance does not depend much on it. For most LMT tasks, there are no more than 20 functions satisfying this requirement. $D_{f,w}>0$ indicates the runtime behavior of function $f$ is unexpected based on our production experience, and $\Delta_{f,w}>M_{f}+kMAD_{f}$ indicates the value of $\Delta_{f,w}$ is significantly larger than most other workers.
As shown in §4.2, the pattern of a worker is only about 30 KB, even for an LMT with 10,000 workers, the analyzer only need to consume 300 MB data to output the final result, which is highly efficient. In our real-world deployment, PerfTracker analyzer is executed on a single CPU core.
Alternatives. To localize abnormal function executions, we also tried existing clustering algorithms, including DBSCAN (Ester et al., 1996), HDBSCAN (McInnes et al., 2017), Gaussian Mixture Model (Bishop and Nasrabadi, 2006) and Mean shift (Comaniciu and Meer, 2002). They do not work well in production because they have at least one of the following limitations: (1) failing to distinguish noises and outliers, which is important in our scenario, or (2) having too many hyper-parameters, reducing their general applicability in practical cases.
5. Implementation
We built PerfTracker with 6K lines of Python code, and deployed PerfTracker as a service in our production GPU clusters. Deploying PerfTracker in production GPU clusters required us to address the following practical challenges.
Optimizations of profiling data generation. PerfTracker relies on Torch profiler to generate a part of raw profiling data (§4). However, Torch profiler introduces overhead to LMT: (1) When the profiling window ends, it requires a long time to prepare and dump the data, and (2) after profiling, it leaves a lot of resources (e.g., hooks for CUDA functions added by CUPTI (cup, 2024)) in the LMT. They may reduce the performance of CUDA kernel execution even after profiling is finished.
PerfTracker addresses the problems by optimizing the Torch profiler. First, the Torch profiler always transfers profiling data to Chrome tracing format, and then dumps the data using Kineto APIs (kin, 2024). Since Kineto supports dumping data with the same format, we remove the redundant and slow format transformation performed by Torch profiler, and directly apply Kineto to dump data. This reduces the time spent on data generation by 33%. Second, PerfTracker calls cuptiFinalize() to clear all remaining resources in LMT after profiling.
Accessing hardware information within the user’s container. In production GPU clusters, an LMT typically runs inside containers; so does PerfTracker daemons (§4.1). However, GPU cluster providers usually restrict the permissions of user containers; therefore, PerfTracker cannot directly access hardware information. PerfTracker utilizes Kubernetes (kub, 2024)’s native “emptyDir” feature to share directories between containers, performing high-frequency hardware sampling through a privileged management container, and placing the data in the shared directory. Therefore, PerfTracker can obtain hardware profiling data within user containers without loosening permissions of user containers.
Cooperating with other hardware monitors. In most production GPU clusters, coarse-grained hardware monitors are deployed on each physical host, using APIs such as DCGM and PCM to collect hardware metrics in real-time for health checks. However, some information (e.g., GPU metrics) can only be subscribed by one process at a time. PerfTracker’s data generator cooperates with the host’s monitoring system via signals (through a shared directory). As PerfTracker’s profiling typically lasts about 20 seconds each time, we do not encounter any conflicts caused by the cooperation.
6. Evaluation in Production
PerfTracker has been deployed in our production GPU clusters with $O(10,000)$ GPUs for three months. During this period, there were in total dozens of LMT performance issues for which we failed to localize the root causes using state-of-the-art techniques described in §2.2 (see Table 2). PerfTracker successfully identified root causes of all of these performance issues, improving the training throughput of these LMTs (the largest one includes 3,072 GPUs) by $\sim$20to $\sim$100%.
We present how we used PerfTracker to diagnose two LMTs with 2,560 and 3,072 GPUs respectively, to show:
  • $\bullet$
    PerfTracker can identify real-world performance issues of both hardware and code-level problems. (§6.1-6.2)
  • $\bullet$
    PerfTracker makes diagnosis in high efficiency, and does not introduce overhead to the routine training of LMT. (§6.3)
  • $\bullet$
    The comparison with the state of the arts. (§6.4)
Table 2. Issues identified by PerfTracker (only includes performance issues not identified by our existing systems)
Category Root cause Proportion of LMT cases
Hardware issues GPU 8.3%
Network 16.7%
Misconfigurations PyTorch 8.3%
Communication 16.7%
Dataloader 8.3%
Low-efficiency code of users 41.7%
6.1. Case 1: Hardware Issues
A text-to-picture LMT job of 2,560 Nvidia H800 GPUs is expected to take 5 seconds per iteration, but is training at 9 seconds per iteration (”expected” and ”original” in Fig. 12).
Figure 12. Iteration time of the LMT in Case 1.
Problem 1: GPU throttling. PerfTracker outputs 300+ workers (located in 40+ hosts) with abnormal execution behaviors on multiple GPU computation kernel functions. Figure 1313(a) shows the example of GEMM (one of the GPU computation kernel functions). GPU computation kernels on these workers have much larger $\beta$ and smaller $\mu$ than most workers, so they have high values of $\Delta_{f,w}$. This indicates that the GPUs of these workers are much slower, so they need a longer time for computation. Interestingly, PerfTracker conducted multiple profiles of this LMT, and in each profile, the slow GPU workers were not consistent, but they are intensively distributed among workers in certain racks, rather than evenly distributed among all workers. This indicates that GPUs from specific deployment batches have intermittent performance issues. Hardware checks revealed recent GPU_NVSMI_HW_THROTTLE alerts on some of these machines, confirming the risk of intermittent GPU throttling. We have contacted the hardware vendor for repairs on these machines.
(a) The value of $\beta$ and $\mu$ (average GPU SM frequency) of GEMM on 2,560 workers.
(b) $\beta$ values of AllGather_RING. There are 2,512 typical points and 48 outliers. We sampled 16 for both of them.
(c) The value of $\mu$ and $\sigma$ (PCIe TX throughput from GPU to NIC) of AllGather_RING on the 48 workers with high $\beta$ value.
Figure 13. Function execution behaviors of GEMM and AllGather_RING in Case 1. Each point represents the runtime behavior pattern of one worker. Outliers are regarded as abnormal function executions with high $\Delta_{f,w}$4.3) and are output by PerfTracker.
Problem 2: NVLink down. PerfTracker also outputs 3 workers with abnormal execution behaviors on multiple collective communication functions. Taking AllGather_RING (one of the collective communication functions) as an example, 48 workers (belonging to three data-parallel groups, where each group has 16 workers) have significantly larger values of $\beta$ than the other 2,512 (Figure 13(b)). Among these 48 workers, three of them have significantly larger $\mu$ than the other 45 (Figure 1313(c)), indicating a higher bandwidth usage of PCIe. Hardware checks revealed NVLink Not Supported (NS) error on these three workers, which means all data flow sent to/from these three workers has to go through PCIe rather than NVLink. Since PCIe is much slower than NVLink, it leads to communication inefficiency.
Fixes. We replaced the problematic hosts with standby hosts and restarted the LMT. Figure 12 shows the iteration time after the replacement (the ”fixed” line).
6.2. Case 2: Code-level Issues
A text-to-video LMT job of 3,072 Nvidia H800 GPUs is expected to take 3.5 seconds per iteration, but is training at 5 seconds per iteration (see the ”original” line in Figure 14).
Figure 14. Iteration time of the LMT in Case 2.
Problem 1: Low throughput in data loader. PerfTracker outputs abnormal function behaviors on the built-in function recv_into of the socket object (a low-level function call in the data loader) on multiple workers. As a GPU-independent function, we expect that the LMT is bottlenecked by it for no more than $1\%$ of the time ($\beta\leq 0.01$, as explained by “expected range” in §4.3). However, as shown in Figure 15(a), there are a lot of workers with a large value of $\beta$ for the recv_into function (so they have high value of $D_{f,w}$). This indicates low efficiency in data loading from the storage.
Problem 2: Inefficient implementation of Python function. Similar to the first problem, PerfTracker outputs abnormal behaviors on the Python function forward with large $\beta$ values, shown in Figure 15(b). This function performs CPU computation and launches GPU kernel functions. A large $\beta$ value indicates the function is bottlenecked by CPU computation, which needs to be optimized.
(a) recv_into of 3,072 workers.
(b) forward of 16 workers. The data of the rest 3,056 are destroyed by the cluster.
Figure 15. CDF of $\beta$ values in two functions.
Problem 3: Asynchronous garbage collection. Apart from the above functions, PerfTracker also reports large $\beta$ values for some other Python functions like gradmode.py:__init__() and _flat_param.py: _get_unflat_views_unaligned. However, these functions are not CPU-intensive, and their executions with large $\beta$ are randomly distributed among all workers. This indicates these workers are executing Python garbage collection asynchronously. Each time a worker is collecting garbage, other workers have to wait for it, leading to GPUs idle and reduced training throughput.
Fixes. For the data loading problem, we suggest that the customer load the input data from our state-of-the-art parallel file system rather than the legacy object storage service, which can significantly reduce the time spent on data loading. For the Python garbage collection problem, we suggest that the customer manually call the garbage collection every 200 iterations. This guarantees all workers collect garbage at the same time, thus avoiding mutual waiting. Implementation optimization of the function forward is not trivial since it contains many CPU computation operations, and we are still working on it. At present, the iteration time of the LMT has reached $\sim$3.6 seconds, as shown in Figure 14.
6.3. System Overhead
Figure 16. PerfTracker’s overhead in timeline view.
Figure 16 illustrates PerfTracker’s overhead. When PerfTracker starts profiling, the LMT is executed with high-precision profiling for a 20-second window (§4.1). When the profiling window ends, LMT will be blocked by data generation for a short period. After that, LMT continues running with no more extra overhead, as the behavior pattern summary is done in a separate process (using a different CPU core), and the root cause localization is located remotely from the worker.
Overhead on production LMTs. We evaluate two LMTs: LMT-A of 2,560 GPUs in §6.1 and LMT-B of 3,072 GPUs in §6.2. Figure 17(a) shows the average iteration time during training with and without profiling. Figure 17(b) shows the duration of data generation, pattern summarization, and localization. The result shows that in production LMT, profiling does not affect the performance of LMT. After that, the localization can be finished within three minutes. In this process, the only overhead on LMT is data generation of $\sim$20 seconds, which is accepted by all of our customers.
(a) LMT iteration time with and without profiling.
(b) Time spent on different components.
(c) Time spent on localization under larger LMT scales.
Figure 17. The overhead introduced by PerfTracker.
Scalability to 1,000,000-GPUs LMTs. In PerfTracker, profiling, data generation and pattern summarization are all distributed in each worker’s container and executed in parallel, so their overhead is independent of the LMT scale. The only overhead proportional to the LMT scale is the centralized root cause localization, which is executed on a single CPU core. We generate simulated runtime behavior patterns as input, to evaluate the overhead of localization for larger-scale LMTs. Figure 17(c) shows that the localization requires only three minutes for a 1,000,000-GPU LMT. Together with the time spent on profiling, data generation, and pattern summarization, PerfTracker can make an end-to-end performance diagnosis for a 1,000,000-GPU LMT (which is larger than any existing LMT we have ever known) within 7 minutes. We cannot make a real evaluation of this, since currently we do not have a customer of such a training scale.
Overhead in different model configurations. Since profiling and data generation are the two components executed in the main process of LMT (Figure 16), Table 3 evaluates them thoroughly in different configurations. In general, the time spent on data generation correlates positively with the number of function executions on a single worker within the profiling window. Increasing the degree of parallelism (e.g., the tensor parallelism, TP) results in more fragmented model training. This fragmentation leads to more function execution events, increasing the data generation time.
Table 3. Overhead in different model configurations
Configurations Overhead
model size tp pp training profiling generate data
(s/iter) (s/iter) (s)
gpt3-7b 1 1 1.371 1.389 13
2 1 1.777 2.002(+12%) 19
gpt3-13b 2 1 2.489 2.485 17
4 1 1.228 1.425(+16%) 26
8 1 1.528 1.707(+11%) 28
gpt3-65b 8 4 1.191 1.202 15
8 8 1.281 1.288 10
In most configurations, profiling does not introduce overhead to training. However, when the model is small and the parallelism parameters are large (e.g., GPT-3 7B with TP=2; GPT-3 13B with TP=4 or 8), there is a major increase in training time. This is because such configurations impose a higher load on the CPU; as profiling requires CPU computation, it causes contention. Fortunately, these impractical configurations are never used in production. Note: PerfTracker is still effective in these configurations and the overhead only happens during the profiling window.
6.4. Comparisons with State of the Arts
We compared PerfTracker with state-of-the-art techniques described in recent papers (Jiang et al., 2024; Deng et al., 2024; Dong et al., 2024), as well as existing profiling tools. Table 4 shows the comparison.
For online monitors, we compare PerfTracker with techniques in MegaScale (Jiang et al., 2024) that records all CUDA events in a timeline (to reveal slow GPU computations) and millisecond- to second-level monitoring of RDMA. Such techniques only trigger alerts or display statistical results; root cause localization (especially for network issues) is usually completed manually. As such techniques do not collect Python function events, they cannot address code-level issues. Similarly, the technique in Minder (Deng et al., 2024) only collects hardware metrics but is unaware of LMT executions. It focuses on LMT crashes, not performance issues. The technique in C4 (Dong et al., 2024) collects runtime information from the communication library, and hence only addresses communication problems.
We also compare PerfTracker with directly applying profilers like Nsight Systems (nsy, 2024) and Torch Profiler (tor, 2024). Nsight provides a Post-Collection Analysis API, which can take profiling data as input for custom analysis. Torch Profiler can dump profiling data as JSON files for offline analysis. Not surprisingly, directly using these profilers online incurs significant overhead in production LMT—even reading fine-grained profiling data is untenable.
Table 4. Comparisons with state-of-the-art approaches
Technique Problems Diagnostic time for 10,000-GPU LMT
C1P1 C1P2 C2P1 C2P2 C2P3
MegaScale (Jiang et al., 2024) online
Minder (Deng et al., 2024) online
C4 (Dong et al., 2024) online
Nsight System ¿1.5 days (offline)
Torch Profiler ¿3.5 days (offline)
PerfTracker 3 min (online)
Table 4 shows the comparison of PerfTracker and the reference techniques in terms of the troubleshooting ability and diagnosis time using the problems in §6.1 and §6.2 (e.g., C1P1 represents Problem 1 of Case 1 in §6.1). Since existing online monitors have incomplete data sources, they cannot troubleshoot many problems. While profilers (Nsight and Torch Profiler) can cover more, loading profiling data of all workers requires days. Since they do not provide troubleshooting logic, we only count the time for data loading in Table 4. Together with troubleshooting, they required even more time.
7. Limitations and Discussions
Scope. PerfTracker does not ensure automatically identifying the root causes of all possible performance issues. (1) While for most performance issues, PerfTracker’s output of the problematic functions and their corresponding runtime behavior patterns is sufficient to determine or closely approximate the root cause, some Python functions and CPU kernels have complex logic. Therefore, even when PerfTracker identifies that these functions have problems (e.g., excessively long execution times), analyzing the root causes still requires manual inspection of their code. (2) As described in §4.3, PerfTracker does not consider any runtime behavior patterns of GPU computation kernels as unexpected. Consequently, if a user manually implements an inefficient GPU computation kernel that runs slowly on all workers, it will not be detected by the distance from expectation or differential distance described in §4.3. PerfTracker requires prior knowledge (e.g., an expected range for this kernel function) to diagnose the issue.
LMTs with non-NVIDIA GPUs or non-PyTroch frameworks. In the clusters where PerfTracker has been deployed, all LMTs (e.g., in Table 2) use Nvidia GPUs. Additionally, PerfTracker utilizes nsys to sample hardware metrics, which is only applicable to Nvidia GPUs. In fact, PerfTracker has also been tested on clusters with non-Nvidia GPUs. In these environments, PerfTracker can still use Torch Profiler to collect Python functions, CPU operations, memory operations, and CUDA kernels, but hardware metrics must be collected using tools specific to other GPU models. Therefore, if the sampling tools provided by these GPUs are less precise than nsys, it may affect the ability to diagnose some hardware-related performance issues. The ability to diagnose code-level problems remains unchanged. In addition, PerfTracker only supports PyTorch at present since PyTorch is the dominant framework. However, the insight and the design of this paper are generic to any frameworks (like TensorFlow (Abadi et al., 2016)).
8. Related Works
Hardware monitoring tools. GPUs can be monitored by a series of tools developed by NVIDIA. NVML (nvm, 2023) is a comprehensive, C-based API library. It provides a standardized interface for users to interact with the GPUs, allowing them to actively sample the status of these devices. DCGM (dcg, 2024) is specifically designed for monitoring GPUs in large data centers. It offers a centralized and scalable monitoring method, making it highly suitable for enterprise environments with extensive GPU deployments. CPUs can be monitored by PCM (pcm, 2024), an application programming interface for monitoring performance and energy metrics of Intel processors. For monitoring NICs (e.g., Mellanox), mstflint (mst, 2024) is developed to access packet or byte-level throughput data.
LMT profiling and analysis. Nsight Systems and Nsight Compute provide comprehensive performance analysis and profiling capabilities. Nsight Systems offers a holistic view of CPU, GPU, and memory usage, while Nsight Compute focuses on detailed kernel performance metrics for CUDA applications, helping developers identify and optimize computational bottlenecks. Torch Profiler is a tool within the PyTorch ecosystem that assists in optimizing PyTorch models by providing detailed insights for each function.
Fast recovery. Fast recovery for LMT is currently achieved through periodic (Microsoft, 2023; T. L. et al., 2023) or incremental (Eisenman et al., 2022) checkpointing of the LMT state to the host memory/disks. Checkfreq (Mohan et al., 2021) mitigates the checkpointing overhead by overlapping computation with the checkpoint process, and employs an adaptive runtime checkpointing frequency through profiling. Gemini (Wang et al., 2023) transfers checkpointing data to multiple hosts to improve robustness. It performs the transmission during the gaps between collective communications. Tanmaey et al. (Gupta et al., 2024) only make a checkpoint after an error occurs, and the faulty rank is recovered from the state of the healthy ranks a minibatch before. This avoids the need to re-initialize the entire training job state. There are other approaches to reduce the checkpointing overhead by asynchronous checkpointing (Nicolae et al., 2020) and partial checkpointing (Maeng et al., 2021).
9. Conclusion
This paper presents PerfTracker, an online diagnosis system for LMT. PerfTracker presents the process of performance degradation detection, behavior pattern summarization, and performance issue localization, which utilize runtime behavior patterns of functions to make diagnosis. We deploy PerfTracker in real-world production GPU clusters and solve sophisticated performance issues for real users.
APPENDIX
Appendix A Examples of LMT executions
Figure 18. The runtime behavior of an MoE training.
Figure 19. The python function behaviors of a forward step in the MoE training.
Figure 20. The kernel function (includes both computation and communication) behaviors of a forward step in the MoE training.
Figure 18 shows the timeline (using https://ui.perfetto.dev/) of high-level Python function call stacks of two adjacent training iterations of one of our production MoE models. Each iteration includes multiple forward and backward phases, and their function runtime behaviors are identical. All workers are executing the same functions with the same behaviors. Due to space limitations, we only present one of them.
Figure 19 and 20 zoom into one of the forward phase in Figure 18. It shows that in the forward phase, low-level functions are also executed repeatedly, and their behaviors (e.g., duration of execution) are highly similar across multiple executions.
Notice that in some configurations, the function runtime behaviors may be slightly different between workers. For example, in pipeline parallel, the workers in the middle pipeline stages require extra send_recvs compared with the first and the last stage. However, function behaviors of workers in the same pipeline stage (in different data parallel groups for example) are still expected to be identical. As a result, the idea of making comparisons of runtime behavior patterns for performance diagnosis still makes sense in these cases.